### v1

Abstract
Superfund sites are locations in the United States with high levels of environmental toxicants, often resulting from industrial activity or improper waste management. Given mounting evidence linking prenatal environmental exposures to adverse birth outcomes, estimating the impact of Superfund remediation is of substantial policy relevance. A widespread approach is to fit a spatial regression, i.e., a linear regression of the outcome (e.g., birth weight) on binary treatment (e.g., indicator for Superfund site remediation) and covariates, along with a spatially structured error term to account for unmeasured spatial confounding. Despite this common practice, it remains unclear to what extent spatial regression models account for unmeasured spatial confounding in finite samples and whether such adjustments can be reformulated within a design-based framework for causal inference. To fill this knowledge gap, we introduce a weighting framework that encompasses three canonical types of spatial regression models: random effects, conditional autoregressive, and Gaussian process models. This framework yields new insights into how spatial regression models build causal contrasts between treated and control units. Specifically, we show that: 1) the spatially autocorrelated error term produces approximate balance on a hidden set of covariates, thereby adjusting for a specific class of unmeasured confounders; and 2) the error covariance structure can be equivalently expressed as regressors in a linear model. We also introduce a new average treatment effect estimator that simultaneously accounts for multiple forms of unmeasured spatial confounding, as well as diagnostics that enhance interpretability. In a study of Superfund remediation, our approach illuminates the role of design-based adjustment for confounding and provides guidance for evaluating environmental interventions in spatial settings.
Keywords: Causal Inference; Spatial Regression; Superfund Sites; Unmeasured Confounding; Weighting Methods
1 Introduction
1.1 Superfund remediation and birth outcomes
The United States contains thousands of hazardous waste sites that potentially expose millions of Americans to toxic pollutants. More than 78 million Americans — 24% of the US population, including 17 million children and 11 million elderly individuals — live within three miles of a hazardous waste site (E.P.A. 2022). Many of these sites are legacies of historical industrial activities, including landfills, manufacturing facilities, mining sites, and processing plants. In 1980, Congress enacted the Comprehensive Environmental Response, Compensation, and Liability Act (CERCLA), granting the Environmental Protection Agency (EPA) authority to identify and remediate these sites, which became known as Superfund sites. Although cleanup efforts have been associated with reductions in congenital anomalies in infants (Currie et al. 2011) and blood lead levels in children (Klemick et al. 2020), remediation remains contentious due to high costs — often tens of millions per site — and disputes over whether the benefits justify the expense (Hamilton and Viscusi 1999).
Our goal is to estimate a causal effect of Superfund site remediation on five birth outcomes while adjusting for a rich set of sociodemographic covariates. A key challenge is unmeasured spatial confounding. Unmeasured spatial confounders are unobserved spatially structured variables that influence both treatment assignment and outcomes, potentially leading to biased effect estimates and invalid confidence intervals (Gilbert et al. 2021; Reich et al. 2021). Examples of unmeasured spatial confounders are community participation and local funding availability, which can shape both cleanup efforts and birth outcomes (Petrie 2006).
1.2 Spatial regression and spatial confounding
To adjust for potential unmeasured spatial confounding, researchers frequently introduce a mean zero spatially autocorrelated error term into a linear regression model and interpret the treatment coefficient as a causal effect, a practice that remains widespread in many disciplines. Throughout this paper, we refer to such models as “spatial regression models” (Anselin 2002). For example, a standard approach is to regress the outcome variable (e.g. birth weight) on a binary treatment variable (e.g. indicator for Superfund site remediation) and measured confounders while incorporating random intercepts at the county or state level. Alternatively, researchers use Gaussian process regression, modeling the correlation between the error terms of two Superfund sites as a function of their geographic proximity, or employ a conditional autoregressive model, which models spatial autocorrelation based on an adjacency or weight structure (Reich et al. 2021).
However, the literature offers conflicting perspectives on whether spatial regression models should be used to adjust for unmeasured spatial confounding. The prevailing view is that spatial regression should not be used to adjust for unmeasured spatial confounding, due to the bias of the treatment coefficient estimate in finite samples (Hanks et al. 2015; Hodges and Reich 2010; Khan and Berrett 2023; Narcisi et al. 2024; Nobre et al. 2021; Paciorek 2010; Page et al. 2017; Schnell and Papadogeorgou 2020). This view also arises from the recognition that a spatial regression model is inherently misspecified when there is unmeasured spatial confounding. Specifically, spatially structured error is typically modeled independently of treatment, under the assumption that it does not capture unmeasured spatial confounding — which, if present, would induce dependence between the error term and the treatment (Hodges and Reich 2010).
Recently, Gilbert et al. (2025) showed that, under certain conditions, spatial regression models with Gaussian process covariance yield consistent treatment effect estimates under unmeasured confounding that is continuous in space. Although the bias of treatment coefficient is non-zero in finite samples, the authors show that it vanishes asymptotically. While this result is of substantial interest, it remains uncertain whether similar consistency guarantees extend to other commonly used spatial regression models — such as random effects models and conditional autoregressive models — or what implications this has for the magnitude of unmeasured spatial confounding bias in finite samples.
1.3 Contribution and outline
The aforementioned literature is subject to limitations. Each approach studies the bias of spatial regression estimators under the restrictive assumption of an outcome model with a linear effect of treatment and without any interactions. Additionally, Gilbert et al. (2025); Nobre et al. (2021); Paciorek (2010); Page et al. (2017) specify a particular form for the error covariance matrix, such as a Gaussian process covariance matrix. The impact of this choice on unmeasured confounding bias is not explored. Dupont et al. (2023); Khan and Berrett (2023); Narcisi et al. (2024) leave the error covariance matrix in a general form, but their bias results — expressed qualitatively in terms of the relationship between the unmeasured confounder and the eigenstructure of the precision matrix — are not straightforward to interpret. Finally, Narcisi et al. (2024); Nobre et al. (2021); Paciorek (2010); Page et al. (2017) assess the bias while imposing strong distributional assumptions on the unmeasured confounder $U$, such as multivariate normality.
In general, the adjustment mechanism by which a spatial regression model accounts for unmeasured spatial confounding remains to be understood, particularly in finite samples, leaving open questions about the hypothetical experiment and the causal contrast it defines in space. In this paper, we provide an exact finite-sample analysis of spatial regression models in binary treatment settings. Our contributions are fourfold. First, we introduce a general weighting framework for causal inference that encompasses three canonical types of spatial regression models: random effects models, conditional autoregressive models, and Gaussian process models. We show that generalized least squares (GLS) estimators derived from spatial regression models are equivalent to minimal dispersion balancing weights estimators, which minimize a measure of weight dispersion subject to covariate balance requirements (Bruns-Smith et al. 2023; Chattopadhyay and Zubizarreta 2023; Kline 2011; Robins et al. 2007). Our equivalence results are the first to show that adjustments made by GLS regression approximately balance the means of a hidden set of additional covariates. The hidden set of additional covariates corresponds to the eigenvectors of the error covariance matrix. This perspective is related to the work of Hodges and Reich (2010), who showed that the spatially correlated error term in a linear model can be represented through additional canonical regressors with random coefficients. Importantly, we further demonstrate that the resulting GLS regression is equivalent to a form of penalized regression, and we provide a mathematical programming representation that characterizes the exact nature of this regression adjustment.
Second, we characterize the class of unmeasured confounders that the three spatial regression models are designed to adjust for. We do so by relating the unmeasured confounding bias due to $U$ to its Moran’s I statistic, which is a measure of spatial autocorrelation of $U$. This mathematical characterization is accompanied by three visual diagnostics to complement traditional weighting diagnostics: a graph of the maximum unmeasured confounding bias in finite sample as a function of the unmeasured confounder’s Moran’s I, illustrative examples of unmeasured confounders that are approximately adjusted for, and a graph of implied weights.
Third, we propose a more general spatial weighting estimator under a causal inference framework that overcomes many of the limitations of spatial regression. Our proposed estimator accounts for multiple forms of unmeasured confounders with spatial autocorrelation, while accommodating nonlinearities and treatment effect heterogeneity by the unmeasured confounder in the outcome model. We derive bounds on the finite-sample bias of our estimator and evaluate its performance in simulation.
Fourth, we apply our proposed spatial weighting estimator to study the impact of Superfund site remediation on five birth outcomes. Among other insights, our approach reveals how the causal contrast of remediated and non-remediated Superfund sites is built across the geography of the United States.
This paper proceeds as follows. Section 2 describes the dataset on Superfund sites and birth outcomes. Section 3 introduces three common examples of spatial regression models, including models with random intercepts at the state level, conditional autoregressive models, and Gaussian process models. Section 4 establishes a general weighting framework that includes GLS estimators derived from spatial regression models as a special case. Section 5 relates the unmeasured confounding bias due to $U$ to its Moran’s I statistic and develops new diagnostics illustrating this result. Section 6 proposes an alternative estimator to spatial regression and derives its finite-sample properties in the presence of unmeasured spatial confounding. Section 7 evaluates the performance of our estimator in simulation and Section 8 uses it to estimate a causal effect of Superfund site remediation on five birth outcomes in New Jersey.
2 Superfund site remediation in the United States
In this study, we examine the impact of Superfund site remediation on five birth outcomes among the sites that were cleaned. The dataset is publicly available and fully reproducible code is accessible at https://github.com/NSAPH-Projects/spatial_regression_weighting.
The dataset includes $n=1583$ Superfund sites in the United States that were listed on the National Priorities List (NPL) as of 2001. The binary treatment variable $Z_{i}\in\{0,1\}$ indicates whether the $i$th Superfund site was remediated and removed from the NPL between 2001 and 2015. Figure 1 plots binary treatment.
Figure 1: Binary Treatment. There were $n=1583$ Superfund sites on the National Priority List on January 1, 2001. During the fifteen-year period between January 1, 2001 and December 31, 2015, $n_{1}=159$ sites were cleaned up and removed from the National Priority List.
We consider ten sociodemographic confounders derived from the 2000 Decennial Census, measured within 2 kilometer buffers around each site. These covariates, originally obtained at the census tract-level and subsequently summarized to the buffer-level, include population density, percentage of individuals identifying as Hispanic, Black, Indigenous, or Asian, median household income, median house value, percentage below the poverty line, percentage with a high school diploma, and the median year of housing construction. We additionally include site score as a covariate — a numerical rating assigned to each site by the EPA during a hazardous risk assessment. Let $\bm{X}_{i}\in\mathbb{R}^{12}$ denote the vector of observed covariates for site $i$, including an intercept.
We consider five birth outcomes measured within $2$ kilometers of site $i$ during the period 2016 to 2020. Importantly, this paper focuses on the design stage of spatial regression and outcome data is not used until Section 8.
Under the Stable Unit Treatment Value Assumption (SUTVA), let $\{Y_{i}(1),Y_{i}(0)\}$ be the potential outcomes under treatment and control, respectively, such that $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$. We assume that the coefficient estimate of treatment obtained from a spatial regression model is interpreted by the researcher, implicitly or explicitly, as the average treatment effect on the treated (ATT). This estimand is defined as $\tau_{ATT}=\mathbb{E}(Y_{i}(1)-Y_{i}(0)|Z_{i}=1).$
3 Three common spatial regression models
In this section, we describe a general form of a spatial regression model and introduce three widely used classes of such models for estimating causal effects in spatial settings. For clarity of notation, let $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{n})^{\top}\in\mathbb{R}^{n\times 12}$ denote the matrix of observed covariates, $\bm{Z}=(Z_{1},\ldots,Z_{n})^{\top}\in\mathbb{R}^{n}$ represent the vector of treatment assignments, and $\bm{Y}=(Y_{1},\ldots,Y_{n})^{\top}\in\mathbb{R}^{n}$ be the vector of outcomes.
The general form of a spatial regression model is
$\bm{Y}=\bm{X}\bm{\beta}+\tau\bm{Z}+\bm{\epsilon},\qquad\bm{\epsilon}\sim\mathcal{N}\left(\bm{0},\;\bm{\Sigma}=\sigma^{2}\bm{I}_{n}+\rho^{2}\bm{S}\right),\qquad\bm{\epsilon}\perp\!\!\!\perp(\bm{X},\bm{Z}),$ (1)
where the covariance matrix of the error terms $\bm{\Sigma}$ includes both an independent and identically distributed term ($\sigma^{2}\bm{I}_{n}$) and a spatially autocorrelated term ($\rho^{2}\bm{S}$). Typically, the $(i,j)$ entry of $\bm{S}$ decreases as the spatial distance between sites $i$ and $j$ increases, reflecting Tobler’s First Law of Geography (Miller 2004). The common assertion that spatial regression cannot adjust for unmeasured spatial confounding stems from the independence assumption, which treats the errors as unrelated to covariates or treatment and thus rules out the presence of unmeasured confounders (Gilbert et al. 2025).
Throughout this work, we assume that the error covariance matrix $\bm{\Sigma}$ is known and specified in advance. Our analysis is motivated by the observational study of Superfund remediation, but it extends to general linear regression models with arbitrary error covariance structures. We focus on the generalized least squares (GLS) estimator for $\tau$, denoted $\widehat{\tau}_{GLS}$, which is given by
$\begin{pmatrix}\widehat{\bm{\beta}}_{GLS}\\ \widehat{\tau}_{GLS}\end{pmatrix}$ $=\bigg{(}\begin{pmatrix}\bm{X}&\bm{Z}\end{pmatrix}^{T}\bm{\Sigma}^{-1}\begin{pmatrix}\bm{X}&\bm{Z}\end{pmatrix}\bigg{)}^{-1}\begin{pmatrix}\bm{X}&\bm{Z}\end{pmatrix}^{T}\bm{\Sigma}^{-1}\bm{Y}.$ (2)
3.1 Random effects models
As a first example, consider a random effects model that incorporates state-level random intercepts. Let $C_{i}$ denote the state indicator for site $i$, taking values $c\in\{1,\ldots,K\}$ with $K=51$, so that $C_{i}=k$ indicates that Superfund site $i$ is in state $k$. The model is
$Y_{i}=\bm{\beta}^{\top}\bm{X}_{i}+\tau Z_{i}+\sum_{k=2}^{K}\gamma_{k}\,I(C_{i}=k)+\xi_{i},\quad\gamma_{k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\rho^{2}),\quad\xi_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2}),\quad(\bm{\gamma},\bm{\xi})\perp\!\!\!\perp(\bm{X},\bm{Z}),$
where $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{K})$, $\bm{\xi}=(\xi_{1},\ldots,\xi_{n})$. In the general form (1), the spatial covariance matrix $\bm{S}$ is block diagonal, with the $k$th block given by $J_{n_{k}}$, an $n_{k}\times n_{k}$ matrix of ones, where $n_{k}=\sum_{i=1}^{n}I(C_{i}=k)$.
As $\rho^{2}\rightarrow\infty$, the model enforces no shrinkage, and the coefficient estimates converge to those obtained from a model with fixed intercepts at the state level. As $\rho^{2}\rightarrow 0$, the model imposes extreme shrinkage, forcing all random intercepts $\gamma_{k}$ to zero. In this case, the coefficient estimates converge to those obtained from a complete-pooling model, in which the data from all states are pooled together (Gelman and Hill 2006). In our implementation of the random effects model in the Superfund data, we fix $\rho^{2}=10$ and $\sigma^{2}=1$.
3.2 Conditional autoregressive models
The second model we consider is the intrinsic conditional autoregressive (ICAR) model,
$Y_{i}=\bm{\beta}^{\top}\bm{X}_{i}+\tau Z_{i}+\phi_{i}+\xi_{i},\quad\bm{\phi}\sim\mathcal{N}\left(\bm{0},\,\rho^{2}(\bm{D}-\bm{A})^{-1}\right),\quad\bm{\xi}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n}),\quad(\bm{\phi},\bm{\xi})\perp\!\!\!\perp(\bm{X},\bm{Z}),$
where $\bm{A}$ is a spatial weight matrix and $\bm{D}$ is a diagonal matrix with entries $D_{ii}=\sum_{j=1}^{n}A_{ij}$. A common choice for $\bm{A}$ is the spatial adjacency matrix, although distance-based matrices can also be used (Earnest et al. 2007). While $\bm{D}-\bm{A}$ is singular and in a Bayesian sense leads to an improper prior for $\bm{\phi}$, this does not affect our theoretical results (see the Supplementary Material). In the general form of (1), we set $\bm{S}=(\bm{D}-\bm{A})^{-1}$. For the Superfund dataset, we implement the ICAR model by fixing $\rho^{2}=10$ and $\sigma^{2}=1$, and defining $A_{ij}=1$ if site $i$ is among the five nearest neighbors of site $j$ or vice versa, and $A_{ij}=0$ otherwise.
3.3 Gaussian process models
The third class is a Gaussian process (GP) model, specified as
$Y_{i}=\bm{\beta}^{\top}\bm{X}_{i}+\tau Z_{i}+\nu_{i}+\xi_{i},\quad\bm{\nu}\sim\text{GP}(\bm{0},\rho^{2}\bm{K}),\quad\bm{\xi}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n}),\quad(\bm{\nu},\bm{\xi})\perp\!\!\!\perp(\bm{X},\bm{Z}),$
where $\bm{K}$ is the kernel matrix of the Gaussian process. In the general form (1), this corresponds to setting $\bm{S}=\bm{K}$.
We apply the GP model to the Superfund dataset by setting $\rho^{2}=10$ and $\sigma^{2}=1$ and using a Matérn correlation with hyperparameters $\kappa=0.2,\phi=\frac{30}{2\sqrt{0.2}}$ and the distance $d_{ij}$ between sites $i$ and $j$ is measured in meters. Specifically, $K_{ij}=\frac{2^{1-\kappa}}{\Gamma(\kappa)}(d_{ij}/\phi)^{\kappa}B_{\kappa}(d_{ij}/\phi)$ where $B_{\kappa}$ is the modified Bessel function of the third kind of order $\kappa$.
4 An encompassing weighting framework
This section presents some of our main results. Our first main result is that the GLS estimator $\widehat{\tau}_{GLS}$ in Equation 2 is a weighting estimator. Our second main result is that $\widehat{\tau}_{GLS}$ is a minimal dispersion balancing weights estimator. In particular, the weights achieve minimal variance while exactly balancing the means of measured covariates and approximately balancing the means of a hidden set of covariates. The latter property characterizes the exact nature of the adjustments done by GLS, clarifying how they emulate key features of a hypothetical experiment and how they differ from other forms of regression adjustments. We note that the results in the following propositions apply to any error covariance matrix $\bm{\Sigma}$; however, we interpret them in the context of spatial regression where $\bm{\Sigma}=\sigma^{2}\bm{I}_{n}+\rho^{2}\bm{S}$ and $\bm{S}$ encodes spatial autocorrelation. The following proposition shows that $\widehat{\tau}_{GLS}$ admits a weighting representation, as it can be re-expressed as a difference of weighted means of the treated and control outcomes.
Proposition 4.1.
The GLS estimator of $\tau$ in Equation 2 can be expressed as
$\widehat{\tau}_{GLS}=\sum_{i:Z_{i}=1}w_{i}Y_{i}-\sum_{i:Z_{i}=0}w_{i}Y_{i}$
with weights $\bm{w}=(w_{1},\ldots,w_{n})$ that admit the following closed-form representation
$\bm{w}$ $=\bm{M}\frac{(\bm{I}_{n}-\bm{\Sigma}^{-1}\bm{X}(\bm{X}^{T}\bm{\Sigma}^{-1}\bm{X})^{-1}\bm{X}^{T})\bm{\Sigma}^{-1}\bm{Z}}{\bm{Z}^{T}\bm{\Sigma}^{-1}(\bm{I}_{n}-\bm{X}(\bm{X}^{T}\bm{\Sigma}^{-1}\bm{X})^{-1}\bm{X}^{T}\bm{\Sigma}^{-1})\bm{Z}},$ (3)
where $\bm{M}$ is the diagonal matrix with $(i,i)$ entry $M_{ii}=2Z_{i}-1$.
Proposition 4.1 shows that the implied unit-level data weights $w_{1},\ldots,w_{n}$ depend only on the treatment indicators, covariates, and error covariance matrix $\bm{\Sigma}$. Therefore, the weights can be obtained as part of the design stage of the study as opposed to the analysis stage. Moreover, the weights can equivalently be obtained by solving a quadratic programming problem.
Proposition 4.2.
Suppose that $\bm{\Sigma}=\sigma^{2}\bm{I}_{n}+\rho^{2}\bm{S}$, where $\bm{S}$ is an $n\times n$ positive semidefinite matrix. Let $\bm{v}_{1},\ldots,\bm{v}_{n}$ be the eigenvectors of $\bm{S}$ with corresponding eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{n}\geq 0$. Consider the following quadratic programming problem:
$\min_{\bm{w}\in\mathbb{R}^{n}}\bigg{\{}\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\bigg{\}}$ (4)
$\text{subject to }\sum_{i:Z_{i}=1}w_{i}=\sum_{i:Z_{i}=0}w_{i}=1\text{ and }\sum_{i:Z_{i}=1}w_{i}\bm{X}_{i}=\sum_{i:Z_{i}=0}w_{i}\bm{X}_{i}.$ (5)
The solution to this problem is the vector of implied weights of $\widehat{\tau}_{GLS}$,
$\bm{w}=\bm{M}\frac{(\bm{I}_{n}-\bm{\Sigma}^{-1}\bm{X}(\bm{X}^{T}\bm{\Sigma}^{-1}\bm{X})^{-1}\bm{X}^{T})\bm{\Sigma}^{-1}\bm{Z}}{\bm{Z}^{T}\bm{\Sigma}^{-1}(\bm{I}_{n}-\bm{X}(\bm{X}^{T}\bm{\Sigma}^{-1}\bm{X})^{-1}\bm{X}^{T}\bm{\Sigma}^{-1})\bm{Z}}.$
Proposition 4.2 demonstrates that $\widehat{\tau}_{GLS}$ belongs to the class of minimal dispersion balancing weights estimators, as it arises from solving a weighting problem that minimizes a measure of weight dispersion (4) subject to covariate balance requirements (5). This result builds on the existing literature that shows numerical equivalences between linear regression and balancing weights (Bruns-Smith et al. 2023; Chattopadhyay and Zubizarreta 2023; Kline 2011; Robins et al. 2007) and also relates to the connection between Gaussian process regression, kernel ridge regression, and weighting (Arbour et al. 2021; Kanagawa et al. 2018). To our knowledge, the reformulation of linear regression models with general error structures as simple weighting estimators offers a new perspective into their finite-sample behavior.
The implied weights of GLS sum to one in each treatment arm and exactly balance the means of the measured covariates included in the regression, similar to both weighted least squares (WLS) and augmented inverse probability weighting (AIPW) (Chattopadhyay and Zubizarreta 2023).
A key distinction of the balancing weights problem for GLS, compared to WLS and AIPW, is its objective function $\{\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\}$. This objective simultaneously minimizes the variance of the weights while penalizing differences in the weighted means of a hidden set of covariates — the eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{n}$ of $\bm{S}$ — between the treated and control groups. These eigenvectors correspond to the canonical regressors identified by Hodges and Reich (2010). When $\sigma^{2}$ is small relative to $\rho^{2}\lambda_{1},\ldots,\rho^{2}\lambda_{n}$, the objective function is dominated by the mean differences of eigenvectors with the largest eigenvalues. As a result, GLS approximately balances the means of these eigenvectors across the treated and control groups — that is, $\sum_{i:Z_{i}=1}w_{i}v_{ki}\approx\sum_{i:Z_{i}=0}w_{i}v_{ki}$ for $\lambda_{k}$ large.
Specifically, the minimal weighting problem of Proposition 4.2 admits an equivalent formulation as a convex quadratic program with a soft-balancing constraint on the sum of the mean eigenvector imbalances. One can instead solve $\min_{\bm{w}}\{\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}\}$ subject to $\sum_{i:Z_{i}=1}w_{i}=1,\sum_{i:Z_{i}=0}w_{i}=1\sum_{i:Z_{i}=1}w_{i}\bm{X}_{i}=\sum_{i:Z_{i}=0}w_{i}\bm{X}_{i},$ and $\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}\leq\Delta(\rho^{2})$ for a specific function $\Delta$. The mapping $\rho^{2}\mapsto\Delta(\rho^{2})$ is derived in Supplementary Section 11.2.
The ratio $\sigma^{2}/\rho^{2}$ parametrizes a balance-dispersion trade-off. A larger ratio favors reducing the variance of the weights, whereas a smaller ratio favors balancing the means of eigenvectors with the largest eigenvalues. This result provides a new perspective on the bias-variance tradeoff in random effects and fixed effects models (see Section 11.2.2), and is of independent interest in the more general context of regression analysis.
Furthermore, the GLS estimator of $\tau$ can be equivalently obtained by augmenting a linear regression of $\bm{Y}$ on $\bm{X}$ and $\bm{Z}$ with the eigenvectors of $\bm{S}$, and applying ridge penalties to their coefficients. Specifically, minimizing the ridge regression loss
$\mathcal{L}(\bm{\beta},\tau,\bm{\gamma})=\sum_{i=1}^{n}\left(Y_{i}-\tau Z_{i}-\bm{\beta}^{\top}\bm{X}_{i}-\sum_{j=1}^{n}\gamma_{j}v_{ij}\right)^{2}+\frac{\sigma^{2}}{\rho^{2}}\sum_{j=1}^{n}\frac{\gamma_{j}^{2}}{\lambda_{j}}$
with respect to $(\bm{\beta},\tau,\bm{\gamma})$ yields $\widehat{\tau}_{\text{ridge}}=\widehat{\tau}_{\text{GLS}}$ (see Supplementary Section 11.3 for details). The ridge penalty on the coefficient for eigenvector $\bm{v}_{k}$ is $\sigma^{2}/(\rho^{2}\lambda_{k})$, so that eigenvectors with the largest eigenvalues are subject to the least shrinkage and effectively enter the linear model as conventional covariates. This perspective also relates to the connection between weighting and kernel ridge regression as discussed in Ben-Michael et al. (2021) and Murray and Feller (2024).
We now interpret this result to explain how spatial regression adjusts for unmeasured spatial confounding. Some intuition can be based on two key observations. First, as discussed above, spatial regression induces approximate mean balance on the eigenvectors of $\bm{S}$ with the largest eigenvalues. Second, these leading eigenvectors correspond to the highest Moran’s I statistics, which is a widely used measure of spatial autocorrelation, where higher values of Moran’s I indicate greater spatial correlation. More formally, the eigenvectors of $\bm{S}$, such as $\bm{v}_{1}=\arg\max_{\bm{x}}\bm{x}^{\top}\bm{S}\bm{x}/\bm{x}^{\top}\bm{x}$, and so forth, maximize the ratio $\bm{x}^{\top}\bm{S}\bm{x}/\bm{x}^{\top}\bm{x}$, which coincides with Moran’s I after centering. Thus, spatial regression implicitly balances the means of a hidden set of unmeasured, spatially autocorrelated covariates. If an unmeasured confounder $\bm{U}$ is highly spatially autocorrelated, it will tend to lie within the span of these eigenvectors, and will therefore be nearly mean-balanced by the spatial regression adjustment. Section 5 formalizes this intuition, providing both a mathematical analysis and visual illustration of the class of unmeasured confounders that spatial regression approximately balances.
5 Understanding spatial regression contrasts across space
Let $\bm{U}=(U_{1},\ldots,U_{n})$ denote an unmeasured confounder, and let $\overline{U}=\frac{1}{n}\sum_{i=1}^{n}U_{i}$ be its sample mean. In this section, we examine how the bias from $\bm{U}$ depends on its spatial autocorrelation structure.
5.1 Causal identification given $U$
Our target estimand is the average treatment effect on the treated (ATT):
$\tau_{\mathrm{ATT}}=\mathbb{E}\big{[}Y_{i}(1)-Y_{i}(0)\mid Z_{i}=1\big{]}.$
To identify this quantity, we first posit the following general assumptions.
Assumption 5.1 (Consistency and no interference).
$Y_{i}=Y_{i}(z)$ when $Z_{i}=z$.
Assumption 5.2 (Positivity of treatment).
$0<\mathbb{P}(Z_{i}=1|X_{i},U_{i})<1$.
Assumption 5.3 (Conditional ignorability given $X,U$).
$\{Y_{i}(1),Y_{i}(0)\}\perp\!\!\!\perp Z_{i}|X_{i},U_{i}$.
From these assumptions, it follows that
$\tau_{\mathrm{ATT}}=\mathbb{E}\Big{[}\mathbb{E}(Y_{i}|Z_{i}=1,X_{i},U_{i})-\mathbb{E}(Y_{i}|Z_{i}=0,X_{i},U_{i})\ \Big{|}\,Z_{i}=1\Big{]}.$
Now we consider the case when $\bm{U}$ is unobserved. In this setting, adjusting for $\bm{X}$ alone generally yields estimators biased by unmeasured confounding. In what follows, we show that when the outcome model is linear and $\bm{U}$ is spatially autocorrelated, spatial regression methods can control the resulting bias in finite samples.
5.2 Spatial autocorrelation and unmeasured confounding bias
We now relate the bias from unmeasured confounding by $\bm{U}$ to the degree of its spatial autocorrelation. Moran’s I statistic is a widely used metric for quantifying spatial autocorrelation. Let $\bm{S}$ be an $n\times n$ positive semidefinite spatial weights matrix with eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{n}\geq 0$. The Moran’s I statistic for $\bm{U}$ relative to $\bm{S}$ is defined as $\mathcal{I}(\bm{U};\bm{S})=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}S_{ij}(U_{i}-\overline{U})(U_{j}-\overline{U})}{\lambda_{1}\sum_{i=1}^{n}(U_{i}-\overline{U})^{2}}$. Although the conventional definition of Moran’s I is $n\lambda_{1}\mathcal{I}(\bm{U};\bm{S})/\sum_{i=1}^{n}\sum_{j=1}^{n}S_{ij}$, the version just presented is normalized to lie between 0 and 1 for ease of interpretation. A value of 1 indicates perfect spatial autocorrelation, while a value of 0 indicates no spatial autocorrelation.
Moran’s I can be interpreted as a scalar summary quantifying how strongly $\bm{U}$ aligns with the eigenvectors of $\bm{S}$ that have the largest eigenvalues. To see this, without loss of generality, assume that $\overline{U}=0$ and $\sum_{i=1}^{n}U_{i}^{2}=1$. Then $\mathcal{I}(\bm{U};S)=\sum_{i=1}^{n}\lambda_{i}(\bm{v}_{i}^{T}\bm{U})^{2}$. In other words, Moran’s I is a weighted sum of the squared coefficients of $\bm{U}$ represented in the eigenbasis of $\bm{S}$, where each coefficient is weighted by the corresponding eigenvalue $\lambda_{i}$. Section 4 suggested that a spatially autocorrelated $\bm{U}$ is largely supported on the leading eigenvectors of $\bm{S}$ with the largest eigenvalues; as a result, $\bm{U}$ is approximately mean-balanced by the weighting procedure of Proposition 4.2. Therefore, spatial regression approximately adjusts for such $\bm{U}$ if the true outcome model is linear. The following proposition formalizes this intuition by relating the bias from unmeasured confounding due to $\bm{U}$ to its Moran’s I statistic.
Proposition 5.1 (Relation between unmeasured confounding bias from $\bm{U}$ and Moran’s I).
Suppose that Assumptions 5.15.3 hold. Further assume that the outcome model is linear, so that $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+\tau Z_{i}+\gamma U_{i}+\epsilon_{i}$, where $\epsilon_{i}$ is an independent and identically distributed error term with $\mathbb{E}(\epsilon_{i})=0$, $\text{Var}(\epsilon_{i})=\sigma^{2}$, and $\epsilon_{i}\perp\!\!\!\perp(\bm{X},\bm{Z},\bm{U})$ for $i=1,\ldots,n$. Let $\bm{\Sigma}=\sigma^{2}\bm{I}_{n}+\rho^{2}\bm{S}$ where $\bm{S}$ is an $n\times n$ positive semidefinite matrix with eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{n}$ and corresponding eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{n}\geq 0$. Then, the bias of $\widehat{\tau}_{GLS}$ for $\tau_{\mathrm{ATT}}=\tau$, conditional on $(\bm{X},\bm{Z},\bm{U})$, is bounded as follows:
$\bigg{|}\mathbb{E}(\widehat{\tau}_{GLS}|\bm{X},\bm{Z},\bm{U})-\tau\bigg{|}$ $=\bigg{|}\gamma(\sum_{i:Z_{i}=1}w_{i}U_{i}-\sum_{i:Z_{i}=0}w_{i}U_{i})\bigg{|}$
$\leq|\gamma|\sqrt{\frac{c_{0}(2\sigma^{2}+\rho^{2}\lambda_{1}+\rho^{2}\lambda_{n})^{2}}{4(\sigma^{2}+\rho^{2}\lambda_{1})(\sigma^{2}+\rho^{2}\lambda_{n})(\sigma^{2}+\rho^{2}\lambda_{1}\mathcal{I}(\bm{U};S))}\sum_{i=1}^{n}(U_{i}-\overline{U})^{2}}$
where $c_{0}=\sigma^{2}\sum_{i=1}^{n}w_{i}^{2}+\rho^{2}\sum_{k=1}^{n}\lambda_{k}(\sum_{i:Z_{i}=1}w_{i}v_{ki}-\sum_{i:Z_{i}=0}w_{i}v_{ki})^{2}$ and $\mathcal{I}(\bm{U};S)$ denotes the Moran’s I statistic of $\bm{U}$ as defined above.
Proposition 5.1 shows that the bias due to unmeasured confounding from $\bm{U}$ tends to decrease as its Moran’s I statistic increases. In other words, spatial regression reduces confounding bias arising from spatially autocorrelated unmeasured confounders. Previous works have arrived at similar results, although through analytical or simulation-based analyses that are less generalizable or interpretable. For example, Paciorek (2010) assume that $\bm{U}$ and $\bm{Z}$ jointly follow a Gaussian process and use the range parameter of the Matérn correlation function to describe spatial autocorrelation, showing that bias decreases as the range of $\bm{U}$ increases. Dupont et al. (2023) quantify spatial smoothness through the eigenvalues of $\bm{S}$ and find that confounding bias is small when $\bm{U}$ aligns with the eigenvectors of $\bm{S}$ corresponding to high eigenvalues. Narcisi et al. (2024) define smoothness in terms of $\bm{U}^{T}\sqrt{\bm{\Sigma}^{-1}}(I-\bm{1}\bm{1}^{T}/n)\sqrt{\bm{\Sigma}^{-1}}\bm{U}$, and show that bias decreases as this measure increases. In contrast, Proposition 5.1 avoids assumptions regarding the distribution or structure of $\bm{U}$ and $\bm{S}$, and makes use of Moran’s I, which is a widely used measure of spatial autocorrelation.
5.3 Visual diagnostics
We now present three diagnostic tools that help clarify the spatial confounding adjustments made by spatial regression models, as well as the composition of the causal contrast they build across the spatial dataset. Our proposed diagnostics build on the weighting diagnostics introduced by Chattopadhyay and Zubizarreta (2023). We apply these diagnostics to the Superfund sites and birth outcomes dataset, examining each of the three spatial regression models described in Section 3.
The first two diagnostics visually characterize the class of unmeasured confounders that spatial regression adjusts for. Figure 2 illustrates the first diagnostic tool, which plots the maximum absolute conditional bias arising from $\bm{U}$ against its Moran’s I statistic and its coefficient in the outcome model. For this diagnostic, we employ nonlinear programming because the analytical bound in Proposition 5.1 can sometimes be quite conservative. For all three spatial regression models, this diagnostic reveals that the maximum absolute conditional bias decreases as the unmeasured confounder’s Moran’s I increases, and increases with the magnitude of $\gamma$, the coefficient of $\bm{U}$ in the outcome model. This pattern empirically corroborates Proposition 5.1.
Figure 2: Maximum absolute conditional bias as a function of Moran’s I and absolute regression coefficient $|\gamma|$.
The second diagnostic, illustrated in Figure 7 of the Supplementary Material, displays three instances of $\bm{U}$ together with their respective Moran’s I statistics and mean imbalances. As expected, confounders characterized by higher Moran’s I tend to exhibit reduced mean imbalance. In particular, the three spatial regression models target unmeasured confounders with different forms of spatial autocorrelation. For the random effects model, spatially autocorrelated unmeasured confounders are nearly constant within states. In contrast, the conditional autoregressive model yields spatially autocorrelated unmeasured confounders that are approximately constant over adjacencies; here, sites $i$ and $j$ are defined as adjacent if site $j$ is one of site $i$’s five nearest neighbors or vice versa. For the Gaussian process model, spatially autocorrelated unmeasured confounders appear to be smooth, continuous functions of spatial coordinates.
Figure 3: Implied weights $w_{1},\ldots,w_{n}$ in three common spatial regression models.
The third diagnostic, shown in Figure 3, displays the implied weights $w_{1},\ldots,w_{n}$ from the three spatial regression models. For the random effects model, the weights are largest in magnitude for Superfund sites located in states with a high proportion of control sites, or for control sites in states with a high proportion of treated sites. For instance, sites in Idaho and Montana, which contain only controls, are assigned weights that are nearly zero. For the conditional autoregressive model, the largest weights correspond to treated sites bordering many control sites, or vice versa. For the Gaussian process model, the largest weights are observed for Superfund sites that are geographically proximate to sites of the opposite treatment group.
In sum, Figure 3 shows that spatial regression produces a weighting estimator in which the treated and control components are weighted to be spatially proximate, according to a definition of proximity that is model-specific. Further discussion of this spatially localized weighting is presented in Supplementary Section 11.5. This phenomenon is akin to distance-adjusted propensity score matching, which pairs treated and control units with similar propensity scores only if they are geographically close (Papadogeorgou et al. 2019) and geographic regression discontinuity designs, which construct treated-control contrasts using observations situated on either side of a geographic boundary separating treatment groups (Keele et al. 2015). However, unlike these methods, spatial regression imposes a very strict requirement of linearity, and if there are any deviations from a simple linear outcome model, its estimates can be subject to large biases. See Supplementary Section 11.6 for examples. In the next section, we present an extension of spatial regression that is capable of addressing nonlinearities and treatment effect heterogeneity in the outcome model.
6 Extending spatial regression estimators for interpretability and robustness
In this section, we extend spatial regression to estimate the average treatment effect on the treated, introducing a weighting estimator that adjusts for multiple forms of unmeasured spatial confounding, relaxes the assumptions of linearity and effect homogeneity, and is constructed to be sample bounded.
Let $X^{\text{aug}}=(\bm{X},\bm{V})\in\mathbb{R}^{n\times(12+J)}$ be the set of measured covariates, augmented with $J$ pre-selected eigenvectors $\bm{V}=(\bm{v}_{1},\ldots,\bm{v}_{J})\in\mathbb{R}^{n\times J}$ of any positive semidefinite matrix $\bm{S}$. Building on Wang and Zubizarreta (2020) and Zubizarreta (2015), we solve the following quadratic programming problem in the control group:
$\min_{\bm{w}\in\mathbb{R}^{n-n_{t}}}\sum_{i:Z_{i}=0}w_{i}^{2}$
$\text{subject to }\big{|}\sum_{i:Z_{i}=0}w_{i}B_{k}(\bm{X}^{\text{aug}}_{i})-\frac{1}{n_{t}}\sum_{i:Z_{i}=1}B_{k}(\bm{X}^{\text{aug}}_{i})\big{|}$ $\leq\delta_{k},k\in 1,\ldots,K,\sum_{i:Z_{i}=0}w_{i}=1,w_{i}\geq 0\text{ }\forall i$
where the functions $B_{k}(\bm{X}^{\text{aug}})$, for $k=1,\ldots,K$, represent a set of pre-specified basis functions, $\delta_{k}$ is the associated imbalance tolerance for the $k$th function, and $n_{t}=\sum_{i=1}^{n}Z_{i}$.
The solution to this optimization problem yields a set of normalized, minimum-variance control weights that approximately balance the basis functions $B(\cdot)=\{B_{1}(\cdot),\ldots,B_{K}(\cdot)\}^{T}$ relative to the treated group. To ensure that the resulting estimator is sample bounded (Robins et al. 2007), we also impose non-negativity constraints on the weights. The spatial weighting estimator is then defined as $\widehat{\tau}_{SW}=\frac{1}{n_{t}}\sum_{i:Z_{i}=1}Y_{i}-\sum_{i:Z_{i}=0}w_{i}Y_{i},$ where $\{w_{i}\}_{Z_{i}=0}$ are the optimized weights obtained from the mathematical program described above.
Our weighting approach requires selecting $J$ augmenting eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{J}$. To reduce bias from the main classes of unmeasured spatial confounding implicitly targeted by the random effects, conditional autoregressive, and Gaussian process models, we include a large number of eigenvectors corresponding to the highest eigenvalues from each model. Unlike methods that adjust for only a single confounder class, $\widehat{\tau}_{SW}$ enables simultaneous adjustment for multiple types of unmeasured confounders, including those constant at the state level, those exhibiting smooth spatial variation, and those with minimal variation across adjacencies. Proposition 11.3 further characterizes the finite-sample bias and variance of the spatial weighting estimator as a function of the Moran’s I statistic of the unmeasured spatial confounder. In the next section, we assess the performance of the spatial weighting estimator in a simulation study.
7 Simulation-based analysis of forms of adjustment
Using the observed covariates $\bm{X}$ and binary treatment $\bm{Z}$ from our Superfund dataset, we create datasets subject to unmeasured spatial confounding. We estimate the ATT using the spatial weighting estimator of Section 6, augmenting the covariate matrix $\bm{X}$ with $J=150$ eigenvectors. The objective of the simulations is twofold. First, we investigate whether the spatial weighting estimator can mitigate confounding bias from three distinct classes of unmeasured confounders with different patterns of spatial autocorrelation. Second, we evaluate its performance while varying the complexity of the outcome model. All simulation code can be found at https://github.com/NSAPH-Projects/spatial_regression_weighting.
For $i=1,\ldots,n=1583,$ we generate unmeasured confounders $\bm{U}$ using three different mechanisms, each inducing a different spatial autocorrelation structure. For all three mechanisms, we generate $\bm{U}$ with a two-stage procedure: 1) we simulate $U_{i}\sim\mathcal{N}(Z_{i},0.1^{2})$ and 2) we set $U_{i}=\sum_{j=1}^{n}W_{ij}U_{j}/\sum_{j=1}^{n}W_{ij}$. The first class of unmeasured confounders (cluster-based) is designed to be constant within states, hence we define $W_{ij}=I(C_{i}=C_{j})$. The second class of unmeasured confounders (adjacency-based) is designed to be smooth across adjacencies, hence we define $W_{ij}=(A^{100})_{ij}$. The third class of unmeasured confounders (distance-based) is designed to vary smoothly with geographical distance, hence we define $W_{ij}=e^{-d_{ij}/50\text{km}}$. Supplementary Section 11.8 provides a plot of one observation of $\bm{U}$ from each class.
We further consider three possible outcome models. The linear outcome model generates outcome as $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+Z_{i}-0.2U_{i}+\mathcal{N}(0,0.1^{2})$. The linear-interaction outcome model generates outcome as $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+Z_{i}-0.5U_{i}+U_{i}Z_{i}+\mathcal{N}(0,0.1^{2})$. The nonlinear-interaction outcome model generates outcome as $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+\tau Z_{i}+Z_{i}\sin(U_{i})-U_{i}^{2}+U_{i}Z_{i}(X_{2i}+1)+0.3Z_{i}X_{1i}^{2}+\mathcal{N}(0,0.1^{2})$. For each of the $9$ data-generating scenarios produced by the three classes of unmeasured confounders and three outcome models, we create $M=500$ datasets of size $n=1583$, with $\bm{X}$ and $\bm{Z}$ drawn directly from our Superfund dataset.
Under each data-generating scenario, we estimate the ATT using six different approaches. The first approach (OLS) fits an ordinary least squares regression of $Y$ on $X$ and $Z$ and extracts the coefficient estimate of $Z$. Approaches two through four employ generalized least squares to fit the random effects (RE), conditional autoregressive (CAR), and Gaussian process (GP) models described in Section 3, likewise extracting the coefficient estimate of $Z$. The fifth approach (spatial coordinates) incorporates the latitude and longitude of each Superfund site as two additional covariates in an augmented inverse propensity score estimator, following the framework proposed by Gilbert et al. (2021). The sixth approach (spatial weighting) implements our proposed methodology from Section 6. In order to adjust for a heterogeneous class of unmeasured confounders with spatial autocorrelation, we augment $\bm{X}$ with $50$ eigenvectors corresponding to the largest eigenvalues from each spatial regression model above, resulting in $J=150$. We use the package sbw for implementation (Zubizarreta 2015).
We evaluate the performance of the six approaches in estimating the ATT through bias and root mean squared error. Figure 4 presents boxplots of the ATT estimates and Supplementary Section 11.8 presents numerical results.
Figure 4: Estimates of the ATT derived from six approaches. OLS, ordinary least squares; RE, random effects; CAR, conditional autoregressive; GP, Gaussian process; SC, spatial coordinates; SW, spatial weighting. The red dotted line represents the true ATT.
We briefly highlight two key observations. First, the results under the linear outcome model, in the first row of the boxplot grid, provide empirical support for our findings in Section 5. Specifically, each spatial regression model yields unbiased estimates for the class of unmeasured confounders it targets: the random effects model under cluster-based confounding; the conditional autoregressive model under adjacency-based confounding; and both the conditional autoregressive and Gaussian process models under distance-based confounding. However, when nonlinearity or treatment effect heterogeneity by $\bm{U}$ is present in the outcome model, all three estimators suffer from bias.
Adjustment for spatial coordinates can reduce bias from unmeasured spatial confounding when the confounder is a measurable, or nearly continuous, function of spatial coordinates and the exposure exhibits non-spatial variation (Gilbert et al. 2021). These assumptions are not explicitly satisfied by any of the confounding mechanisms considered, and the spatial coordinates estimator exhibits substantial bias across most data-generating scenarios.
In contrast, our proposed spatial weighting estimator remains approximately unbiased across all nine data-generating scenarios and — in terms of bias — outperforms each competing model in seven of the nine scenarios, although sometimes exhibits larger uncertainty (see Supplementary Section 11.8). These simulation results demonstrate that the spatial weighting estimator can flexibly accommodate complex nonlinearities and treatment effect heterogeneity, while simultaneously adjusting for unmeasured confounders that are constant within states, smooth across adjacencies, or vary continuously over space.
8 Effects of Superfund remediation on birth outcomes
We now apply our proposed spatial weighting approach to estimate the effect of Superfund site remediation during 2001–2015 on five birth outcomes measured during 2016–2020 in New Jersey $(n=139)$. All five birth outcome variables were obtained at the municipal-level from the New Jersey State Health Assessment and subsequently aggregated to the 2 km buffer level around each Superfund site.
The first outcome is the percent of low birth weight (birth weight $\leq 2500$ grams) births; the second outcome is average birth weight in pounds; the third outcome is infant mortality (death $\leq 1$ year of age) rate; the fourth outcome is preterm birth rate (gestational age <37<37 weeks); and the fifth outcome is average gestational age. These publicly available, population-wide data cover all births, regardless of insurance status or hospital affiliation, at a fine spatial resolution (across $564$ municipalities). New Jersey contains the highest number of Superfund sites out of all US states. Our target of estimation is the causal risk ratio, defined as $\mathbb{E}(Y(1)|Z=1)/\mathbb{E}(Y(0)|Z=1)=\mathbb{E}(Y|Z=1)/(\mathbb{E}(Y|Z=1)-\tau_{ATT})$. Standard errors are derived from bootstrapping with $500$ replications. Figure 10 plots the estimated causal risk ratio as the number of hidden covariates ($2J$) incorporated into the spatial weighting method increases. For $J\in\{1,\ldots,20\}$, we append a total of $J$ highest-eigenvalue eigenvectors of $\bm{S}^{\text{CAR}}$ and $J$ highest-eigenvalue eigenvectors of $\bm{S}^{\text{GP}}$ to the design matrix. The point estimates, although accompanied with large uncertainty, tend to diverge from the null until $J=9$ and then plateau. Since the stabilization occurs at $9$, we adopt $J=9$ for a total of $18$ hidden covariates in the final specification of the spatial weighting estimator.
Table 1 reports estimated causal risk ratios for each of the five birth outcomes under each of the six modeling approaches. Point estimates are broadly similar across the six models, suggesting that unmeasured spatial confounding bias may be modest. The spatial weighting approach yields the lowest estimates for low birth weight and infant mortality — 0.93 (95 % CI 0.86, 1.02) and 0.67 (95 % CI 0.50, 1.01), respectively — with smaller uncertainty than the other four spatial models.
Superfund site remediation may improve birth outcomes by reducing maternal exposure to hazardous substances such as heavy metals, solvents, and air- or waterborne contaminants — such as arsenic, lead, and PFAS — known to affect fetal development (Henn et al. 2016; Keeler et al. 2023). Cleanup efforts may also lead to secondary improvements in neighborhood conditions (e.g., reduced stress, improved housing or infrastructure, neighborhood investment, increased housing prices) that may indirectly improve birth outcome measures (Currie 2011; Kiel and Zabel 2001; Noonan et al. 2007).
Our findings are somewhat consistent with prior research. Currie et al. (2011) report that remediation may reduce the incidence of congenital anomalies by 20–25%, and also find a modest reduction in preterm birth, although no significant effects on low birth weight or infant mortality. In related work, Currie (2011) documents a demographic shift in the population of mothers residing near cleaned sites, with increases in the proportion of White and college-educated mothers, perhaps due to housing price appreciation. Such demographic shifts may contribute to the estimated effects observed in our study. However, the existing literature is limited; most prior research has focused on associations between residential proximity to Superfund sites and birth outcomes, rather than on estimating the causal impact of site remediation (Brender et al. 2011; Henn et al. 2016; Kihal-Talantikite et al. 2017; Langlois et al. 2009).
Rate of Low Birth Weight Average Birth Weight Infant Mortality Rate Preterm Birth Rate Average Gestational Age
OLS 0.96 (0.90, 1.03) 1.00 (1.00, 1.01) 0.83 (0.63, 1.21) 0.96 (0.90, 1.03) 1.00 (1.00, 1.00)
RE 0.96 (0.87, 1.07) 1.00 (1.00, 1.01) 0.80 (0.56, 1.42) 0.96 (0.87, 1.06) 1.00 (1.00, 1.00)
CAR 0.94 (0.85, 1.05) 1.00 (0.99, 1.01) 0.78 (0.54, 1.36) 0.96 (0.88, 1.06) 1.00 (1.00, 1.00)
GP 0.96 (0.87, 1.07) 1.00 (0.99, 1.01) 0.81 (0.57, 1.37) 0.96 (0.88, 1.07) 1.00 (1.00, 1.00)
SC 0.96 (0.87, 1.06) 1.00 (0.99, 1.01) 0.81 (0.59, 1.28) 0.96 (0.88, 1.05) 1.00 (1.00, 1.00)
SW 0.93 (0.86, 1.02) 1.00 (0.99, 1.00) 0.67 (0.50, 1.01) 0.96 (0.88, 1.07) 1.00 (1.00, 1.00)
Table 1: Estimates of the causal risk ratio and 95% confidence intervals from six methods, assessing the impact of Superfund site remediation (2001–2015) on five birth outcomes (2016–2020) within 2 kilometers of Superfund sites in New Jersey. We define the causal risk ratio as $\mathbb{E}(Y(1)|Z=1)/\mathbb{E}(Y(0)|Z=1)$. The five birth outcomes are: 1) average birth weight (pounds), 2) gestational age (weeks), 3) infant mortality rate (%), 4) rate of low birth weight births (%), and 5) preterm birth rate (%). OLS, ordinary least squares; RE, random effects; CAR, conditional autoregressive; GP, Gaussian process; SC, spatial coordinates; SW, spatial weighting.
Figure 5 presents diagnostic plots summarizing the spatial weighting procedure: panel (a) shows covariate balance, including the 18 hidden spatial covariates; panel (b) displays a histogram of the weights; panel (c) maps the weights in geographic space; and panel (d) plots each of the hidden spatial covariates. Approximate balance on the 18 hidden spatial covariates yields a weighted contrast between treated and control units that are geographically close. Although the effective sample size is only $(\sum_{i=1}^{n}w_{i}^{2})^{-1}=45.9$, limiting statistical power, our analysis demonstrates the utility of spatial weighting in a challenging setting. Specifically, it highlights the method’s interpretability, its ability to incorporate spatial information via hidden spatial covariates, and its potential for transparent diagnostic evaluation.
Figure 5:
Control Treated
Covariate Before weighting After weighting
population density 2504.165 3697.645 3697.645
proportion Hispanic 0.097 0.075 0.075
proportion Black 0.110 0.115 0.115
proportion Indigenous 0.006 0.006 0.006
proportion Asian 0.052 0.046 0.046
median household income 59855.855 62735.665 62735.665
median house value 164254.979 183720.081 183720.081
proportion in poverty 0.073 0.056 0.056
proportion graduated HS 0.817 0.852 0.852
median year built 1966.374 1962.972 1962.972
Site Score 45.654 42.085 42.085
$\bm{v}_{1}^{\text{CAR}}$ 0.003 -0.019 -0.017
$\bm{v}_{2}^{\text{CAR}}$ -0.003 0.018 0.021
$\bm{v}_{3}^{\text{CAR}}$ 0.002 -0.009 -0.012
$\bm{v}_{4}^{\text{CAR}}$ 0.003 -0.015 -0.018
$\bm{v}_{5}^{\text{CAR}}$ -0.003 0.015 0.018
$\bm{v}_{6}^{\text{CAR}}$ 0.003 -0.015 -0.020
$\bm{v}_{7}^{\text{CAR}}$ -0.004 0.023 0.024
$\bm{v}_{8}^{\text{CAR}}$ 0.001 -0.004 -0.010
$\bm{v}_{9}^{\text{CAR}}$ 0.004 -0.016 -0.024
$\bm{v}_{1}^{\text{GP}}$ -0.085 -0.085 -0.085
$\bm{v}_{2}^{\text{GP}}$ 0.003 -0.022 -0.021
$\bm{v}_{3}^{\text{GP}}$ 0.000 -0.002 0.002
$\bm{v}_{4}^{\text{GP}}$ 0.001 -0.004 -0.004
$\bm{v}_{5}^{\text{GP}}$ 0.001 -0.011 -0.007
$\bm{v}_{6}^{\text{GP}}$ -0.005 0.030 0.037
$\bm{v}_{7}^{\text{GP}}$ 0.006 -0.030 -0.041
$\bm{v}_{8}^{\text{GP}}$ 0.001 0.003 -0.005
$\bm{v}_{9}^{\text{GP}}$ -0.002 0.015 0.013
(a) Balance and representativeness.
(a) Weight dispersion.
(b) Spatial weights $w_{1},\ldots,w_{n}$.
(c) $18$ hidden covariates — the nine highest-eigenvalue eigenvectors from both $\bm{S}^{\text{CAR}}$ and $\bm{S}^{\text{GP}}$ .
9 Practical considerations
We recommend that researchers examine the components of the spatial contrast to understand how information is shared across space and assess whether this aligns with subject-matter knowledge or intuition. To support this, we conclude by describing how borrowing information across space involves navigating a set of tradeoffs.
First, borrowing strength across space reflects a fundamental bias–variance tradeoff. At one extreme, imposing tight balancing constraints on the hidden spatial covariates restricts comparisons to the nearest treated and control units, so information is borrowed only locally. At the other extreme, relaxing these constraints entirely disregards the spatial adjustment performed by the hidden covariates and borrows information globally. The former approach reduces bias from the targeted form of unmeasured spatial confounding at the cost of greater variance, whereas the latter approach lowers variance but is more susceptible to unmeasured spatial confounding bias. This result mirrors the analytical result in 11.2.1 and aligns with the intuition behind donor pool selection in synthetic control analyses (Abadie 2021). We therefore recommend that investigators (i) carefully examine the distribution of the weights in space and (ii) assess the sensitivity of point estimates and standard errors across a plausible range of balancing thresholds for the hidden covariates.
Second, borrowing strength across space introduces a tension between adjusting for unmeasured spatial confounding bias and avoiding bias from spatial interference. When interference is present, using the most proximal control units as comparisons may be inappropriate, as these units could be affected by the treatment received by nearby treated units. While we do not address spatial interference in the current work, previous studies have proposed selecting control units beyond a spatial buffer, so they are not exposed to the treatment but still have similar values of the unmeasured spatial confounder (Butts 2021). We recommend that, in settings where interference is plausible, researchers consider such buffered comparisons to reduce potential contamination.
A third tradeoff introduced by borrowing strength across space relates to the positivity assumption and the target population. When borrowing strength only locally, some treated units may be too distant from any control units to permit meaningful comparisons. In such cases, we argue that these treated units should be excluded from the target population, as there is no justification for assuming the existence of control units with similar values of the unmeasured confounder (a violation of positivity). Thus, imposing strict balancing constraints on the hidden covariates may require us to redefine the target population. In light of this, we recommend that researchers carefully consider the choice of target population before proceeding with any analysis. All of these comments are in the spirit of matching and weighting methods for observational studies (Zubizarreta et al. 2023). Embedding spatial-contrast considerations into routine practice will strengthen the credibility and interpretability of spatial causal analyses.
10 Discussion
In this paper, we unify three types of spatial regression through a weighting framework for causal inference. Our exact, finite-sample analysis reveals that spatial regression models adjust for unmeasured spatial confounding by approximately balancing a hidden set of spatially autocorrelated covariates determined by the spatial error structure. Furthermore, we show that the confounding bias due to an unmeasured confounder $U$ is small whenever $U$ has high Moran’s I. Our mathematical characterization is accompanied by three visual diagnostics to better understand this result. Finally, we propose a more general spatial weighting estimator that accommodates nonlinearities and treatment effect heterogeneity in the outcome model and simultaneously adjusts for unmeasured confounders with different forms of spatial autocorrelation.
From a regression modeling perspective, our main result makes explicit how the assumed error structure in spatial regression adjusts for unmeasured confounders. Specifically, we show that incorporating a spatial error term is equivalent to augmenting the linear model with additional auxiliary covariates (the eigenvectors of the covariance matrix) whose coefficients are penalized, thereby inducing an approximate mean balance on these covariates. In this sense, “modeling the error” and “softly balancing eigenvectors of the covariance matrix” are two sides of the same coin. More broadly, under a linear model, the error covariance structure can be re-expressed as additional covariate structure with appropriate regularization, making these implicit adjustments explicit and more interpretable.
We highlight two avenues for future work. First, although we provide an exact, finite-sample analysis that informs large-sample considerations, the development of a comprehensive asymptotic theory remains open. Gilbert et al. (2025) established consistency for Gaussian process regression, and Datta and Stein (2025) derived similar results for settings where the exposure, outcome, and unmeasured confounder are generated by Gaussian random fields, both under infill asymptotics. However, extending this theory to random effects or conditional autoregressive models presents additional challenges that warrant further investigation. As $n$ grows, new units may join existing clusters and form new adjacencies, so pre-existing entries of the covariance matrix change with sample size. Accounting for this evolving geometry is essential to any asymptotic treatment of these models.
Second, practitioners typically select the spatial hyperparameter $\rho^{2}$ in a Bayesian framework using outcome information. Here we treated $\rho^{2}$ and $\bm{\Sigma}$ as fixed and known a priori. Our results elucidate how any $\bm{\Sigma}$ — known or unknown, correctly or incorrectly specified — adjusts for confounders in finite samples. Supplementary Section 11.2.1 additionally establishes the role of $\rho^{2}$ in a balance-dispersion (hence bias-variance) tradeoff. Further investigation is needed to understand how outcome-driven hyperparameter selection, including approaches that leverage multiple outcomes (DellaVigna et al. 2025), can inform confounding adjustment in a frequentist framework. This includes the selection of $\rho^{2}$ and $\bm{\Sigma}$, or equivalently, eigenvectors and balance tolerances.